OpenStudy (anonymous):

I need help finding the flux through an ellipsoid (the surface) where F= and x^2+y^2+4z^2=16. How do I go about doing that? I have to evaluate it using the divergence theorem and through a method that involves a sphere through the origin. please help!

OpenStudy (anonymous):

Use the divergence theorem http://en.wikipedia.org/wiki/Divergence_theorem

OpenStudy (anonymous):

Div(F)=2 x + 4 y + 6 z The flux through the skin of the ellipsoid is equal to the integral of the Div(F) over the meat inside that skin. \[ \int _{-4}^4\int _{-\sqrt{16-x^2}}^{\sqrt{16 -x^2}}\int _{-\frac{1}{2} \sqrt{-x^2-y^2+16}}^{\frac{ 1}{2} \sqrt{-x^2-y^2+16}}(2 x+4 y+6 z)dzdydx \] Since Div(F) is odd in x, in y and in z and the region is symmetric with respect to x, y and z, the above integral is equal to zero.